The length of a rectangle is 6 inches more than its width. The area of the rectangle is 91 square inches. Find the dimensions of the rectangle.

Answer by checkley79(3341)

AREA=LW
L=W+6
91=(W+6)W
91=W^2+6W
W^2+6W-91=0
(W-7)(W+13)=0
W-7=0
W=7 ANS.
W+13=0
W=-13 ANS.